------------------------------------------------------------ Date: Mon Aug 14 07:38:08 2000, Shift: 00:00-08:00 Supervisor: RW, with: EJK Log book#4 pages 84 - 86 Shift Summary: ------------------------------------------------------------ 1. Beam was available from 1:40am to 2:30am and 5:00 to 6:00am. Beam intensities were 6.3E9 in Blue and 4.6E9 in Yellow for the first one and 7.5e9 in Blue and 3.5e9 in Yellow in the second time. At 6:00am, Cryogemic controll problem occurs and the beam is lost. 2. Run2248 was finished. MRS at 90 deg. and 4KGauss (700A) field. FFS is the same as before. 25356 collisions were contained. 3.At the second beam run, ZDC HV's were found off. We tried to turn ON from the RHIC terminal, but it didnot work. On the ZCAL HV menu, nom.status and status were found all off. We clicked the nom.status to ON position (this is the regular process to turn ZDC on), but the Vmeas dis not come up. This ZDC HV unit should be taken care of by Main Controll room. No body on shift can handle this problem. We decided to wait until morning. It was 5:30am. T2b(A1.6) could not hold the HV(1550V). T2b stayed off during the run2248. Anyway T2 readout is still broken. ------------------------------------------------------------
This archive was generated by hypermail 2b29 : Mon Aug 14 2000 - 07:42:56 EDT